OpenStudy (anonymous):

What is the simplified form of the expression square root 1 over 121???

OpenStudy (mathstudent55):

This? \( \sqrt{ \dfrac{1}{121} } \)

OpenStudy (anonymous):



OpenStudy (mathstudent55):

\( \sqrt{ \dfrac{1}{121} } \) \( = \dfrac{ \sqrt{1}} {\sqrt{121} } \)

OpenStudy (anonymous):

i don't think you want to write \(\sqrt{1}\)

OpenStudy (mathstudent55):

Notice that the numerator and denominator are perfect squares so you can take the square roots.

OpenStudy (calculusxy):

Just find the square root of 121 and with the square root put it as the denominator and 1 as the numerator.

OpenStudy (anonymous):